A consumer's utility function is given by the

expression: = (0.6X" + 0.4Y°'12.

·    Determine the marginal utility functions for each commodity. Does marginal utility decrease when consumption increases?

·    Assuming that the price of good is Rs 15 and the price of Y is Rs 6, write the equation of the budget line and plot it when income is Rs 450. What is its slope? What does it indicate?

·    Calculate the marginal rate of substitution of Y for X and interpret its economic meaning. Write the equation showing the consumer's equilibrium condition.

·    Obtain the equilibrium values of X and Y.

·    Find the expressions for change in MU), due to increase in Y and change in MU, due to increase in X.

1
Expert's answer
2021-02-17T12:34:44-0500
"U=12x^{0.6}y^{0.4}"

"MU_x=7.2(\frac {y}{x})^{0.4}"

"MU_y=4.8(\frac {x}{y})^{0.6}"


"p_x=15"


"p_y=6"

"15x+6y=240"


"y=-2,5x+75"


"\frac {MU_x}{p_x}=\frac {MU_y}{p_y}"


"x=0.6y"

"y=-2.5\times0.6y+75"


"y=30"


"x=18"

"\frac {MU_y}{MU_x}=\frac {2}{3}\times(\frac {18}{30})=0.4"